# A Gaussian Process View on Observation Noise and Initialization in Wide Neural Networks

### v2

@@ Abstract
Performing gradient descent in a wide neural network is equivalent to computing the posterior mean of a Gaussian Process with the Neural Tangent Kernel (NTK-GP), for a specific prior mean and with zero observation noise. However, existing formulations have two limitations: (i) the NTK-GP assumes noiseless targets, leading to misspecification on noisy data; (ii) the equivalence does not extend to arbitrary prior means, which are essential for well-specified models. To address (i), we introduce a regularizer into the training objective, showing its correspondence to incorporating observation noise in the NTK-GP. To address (ii), we propose a shifted network that enables arbitrary prior means and allows obtaining the posterior mean with gradient descent on a single network, without ensembling or kernel inversion. We validate our results with experiments across datasets and architectures, showing that this approach removes key obstacles to the practical use of NTK-GP equivalence in applied Gaussian process modeling.
@@ 1 Introduction
The connection between wide neural networks and Gaussian Processes (GPs) via the Neural Tangent Kernel (NTK) (jacot2018neural) provides a powerful framework for understanding the training dynamics of deep learning. The NTK describes how predictions of a neural network evolve under gradient-based optimization in the infinite-width limit and has primarily been used to analyze generalization, convergence, and expressivity of neural networks from a theoretical perspective. While this line of work has yielded insights into training dynamics, its practical utility for GP inference remains limited. In particular, the standard NTK-GP correspondence, as established by lee2019wide, connects the output of a wide neural network trained with gradient flow on mean squared error (MSE) to the posterior mean of a GP with the NTK as its kernel–under restrictive conditions. Specifically, this result assumes (i) zero observation noise, leading to model misspecification when learning from real-world data, and (ii) a fixed prior mean given by a randomly initialized network, which is uninformative and not easily controllable. These constraints make the GP equivalence difficult to leverage in practice, as they limit the model’s ability to incorporate uncertainty and prior knowledge, two key concepts of principled Bayesian inference. In this work, we address these limitations and extend the theory in a way that enables practical posterior mean inference with NTK-GPs using a single neural network training run.
@@
Concerning the first limitation, observation — or aleatoric — noise is a fundamental component of Gaussian Processes and other probabilistic models (williams2006gaussian; saez2024neural; ordonezpartially) that represents observation uncertainty. Measurements in data are subject to precision limits, and labels can carry errors due to annotation noise or ambiguities (kendall2017uncertainties). By capturing this variability, aleatoric noise enables models to produce predictions that reflect the randomness of the observed process (williams1998bayesian; bergna2024uncertainty). In GPs, this is achieved through a variance term that accounts for label noise, contributing to robust and well-calibrated predictions.
@@
hu2020simpleeffectiveregularizationmethods tried to address the problem of observation noise by introducing a regularizer penalizing the distance between network parameters at time $t$, $\theta_{t}$, and their initialization, $\theta_{0}$, demonstrating its efficacy in noisy settings. Their analysis, however, relied on the NTK framework and assumed that the network remains in a linear regime throughout training—a critical assumption that was left unproven. This oversight is significant, as the regularizer modifies the training dynamics and invalidates the application of previous results in lee2019wide. Other recent works have used the regularizer proposed by hu2020simpleeffectiveregularizationmethods to improve generalization and training stability. For instance, nitanda2020optimal employed it to constrain networks near their initialization under the NTK regime, while suh2021non extended this to deeper networks for recovering ground-truth functions from noisy data. he2020bayesian explored the regularizer’s role in Bayesian deep ensembles.
@@ 2 Preliminaries
The choice of parametrization in neural networks determines how signals propagate and how gradients scale with the network width. Let $\phi:\mathbb{R}\to\mathbb{R}$ denote the activation function. [^1] We assume that $\phi$ and $\phi^{\prime}$ are Lipschitz. In the standard parametrization, the layer outputs are defined (for $l=0,\ldots,L$) as:
@@
where $W^{l+1}\in\mathbb{R}^{n_{l+1}\times n_{l}}$ and $b^{l+1}\in\mathbb{R}^{n_{l+1}}$ are the weights and biases, respectively. The parameter vector $\theta\in\mathbb{R}^{p}$ is defined by stacking them. Here, $n_{0}:=d$ is the dimension of the input $x^{0}=x$, and $n_{L+1}:=k$ is the dimension of the output $f(x,\theta):=h^{L+1}$. For simplicity, we only consider $k=1$. The initial weights $\theta_{0}$ are i.i.d. $W^{l}_{0,ij}\sim\mathcal{N}(0,\frac{\sigma_{w,l}^{2}}{n_{l}})$, and the biases as $b^{l}_{0,i}\sim\mathcal{N}(0,\sigma_{b,l}^{2})$ [^2] Note, that $W^{1}x^{0}$ sums over $d$ terms, and for $l\geq 1$, $W^{l+1}x^{l}$ sums over $n$ terms.. Under this parametrization, the norm of the Jacobian diverges for width $n_{l}\to\infty$. Without loss of generality, we will assume $n_{1}=\ldots=n_{L}=:n$.
@@
The NTK characterizes the evolution of wide neural network predictions as a linear model in function space. Given a neural network $f(x,\theta)$ parameterized by $\theta\in\mathbb{R}^{p}$, define the Jacobian in any $N$ points $x$ as $J(x,\theta):=\frac{\partial f(x,\theta)}{\partial\theta}\in\mathbb{R}^{N\times p}$. Under NTK parametrization, the empirical NTK at two sets of inputs $x^{\prime}$ and $x$ is defined as:
$\hat{\mathbf{\Theta}}_{x^{\prime},x}:=J(x^{\prime},\theta_{0})J(x,\theta_{0})^{\top}\in\mathbb{R}^{N^{\prime}\times N}.$ (3)
Here, $x^{\prime}$ denotes the set of input points at which predictions are made. Note, that the empirical NTK depends on the randomly initialized $\theta_{0}$. Interestingly, as shown by jacot2018neural, in the infinite width limit, the empirical NTK converges to a deterministic (i.e., independent of $\theta_{0}$) kernel $\mathbf{\Theta}$ (the analytical NTK) at initialization and remains constant during training (with unregularized gradient flow). A GP defined by the analytical NTK is referred to as the NTK-GP.
@@ 3 Motivation
Our work builds on a well-known result (lee2019wide) which shows that, for any sufficiently [^3] In our proofs, we make the notion of ‘sufficiently wide’ formal. wide neural network initialized to parameters $\theta_{0}$, performing gradient descent on the loss
$\frac{1}{2}\sum_{i=1}^{N}(f(\mathbf{x}_{i},\theta)-\mathbf{y}_{i})^{2}$ (4)
@@
We answer this question in the affirmative. Specifically, we build up the theory for supporting $\beta>0$ in Section 4 (Theorems 4.1 and 4.3) and then add the theory for supporting arbitrary prior means in Section 5 (Theorem 5.1).
@@ 4 Observation Noise through Regularized Gradient Descent
The introduction of weight-space regularization not only impacts the generalization properties of wide neural networks (hu2020simpleeffectiveregularizationmethods), but also fundamentally alters their training dynamics. In the results proposed by lee2019wide —where the unregularized gradient flow is studied— these dynamics are crucial to understanding how networks converge to solutions that (in the mean posterior sense) align with Bayesian inference. In this section, we analyze the training dynamics of wide neural networks under regularized gradient flow, with a particular focus on the consequences of regularization on the linearization of the network during training. We demonstrate that the training trajectory stays arbitrarily close to its linearized counterpart, providing both theoretical justification and insights into how regularization modifies the gradient flow.
Define $f(\theta):=f(\mathbf{x},\theta)$, $g(\theta):=f(\mathbf{x},\theta)-\mathbf{y}\in\mathbb{R}^{N}$, $J(\theta):=J(\mathbf{x},\theta)\in\mathbb{R}^{N\times p}$ in the training points. [^4] We assume that $\mathbf{x}_{i}\neq\mathbf{x}_{j}$ for $i\neq j$. We will consider (for NTK parametrization [^5] See Appendix E.2 for standard parametrization.) the regularized training loss
@@
We derive this closed-form solution (for gradient flow and gradient descent) in Appendix D.
Moreover, if we consider the network initialization as a random variable, the trained linearized network converges to a normal distribution for layer width $n\to\infty$. Its mean corresponds to the posterior mean of the NTK-GP with prior mean 0. However, similarly to lee2019wide; he2020bayesian, the covariance cannot be interpreted as the posterior covariance of an NTK-GP. See Appendix I for the exact formulas.
In the remainder of this section, we show that, with high probability over initialization, the training dynamics of the network (at any test point $\mathbf{x^{\prime}}$) closely follow the output of the linearized network for large enough hidden layer width $n$. First, we leverage the fact that the Jacobian is locally bounded and locally Lipschitz, with a Lipschitz constant of $O\left(\frac{1}{\sqrt{n}}\right)$ (see Appendix A for details). Next, we prove that the distance between the parameters and their initialization $\lVert\theta_{t}-\theta_{0}\rVert_{2}$ is O(1)O(1) [^8] On average, this implies that the distance of individual parameters barely changes.. This step is particularly challenging because proofs for $\beta=0$ (in Eq. (6)) rely on the MSE $\lVert g(\theta_{t})\rVert_{2}$ decaying exponentially to 0, which does not hold for $\beta>0$. As this is not the case for $\beta>0$, we instead show that $\lVert\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})\rVert_{2}$ decays to 0 exponentially fast.
@@ 4.1 Exponential Decay of the Regularized Gradient and Parameter Proximity to Initialization
Building upon the established Lipschitz continuity of the Jacobian (in Appendix A), we analyze the effect of regularized gradient flow on the norm of the gradient and the parameters’ closeness to their initialization. We begin by demonstrating that the norm of the gradient of the regularized loss decays exponentially over time. This result directly implies that the distance between the parameters and their initialization is bounded by a constant, i.e. $\lVert\theta_{t}-\theta_{0}\rVert_{2}=O(1)$ [^9] The bounding constant is independent of the outcome in the high probability event for which this holds.. Furthermore, leveraging the Lipschitzness of the Jacobian established in Lemma A.1, we show that the Jacobian remains close to its initial value, with deviations scaling as $O\left(\frac{1}{\sqrt{n}}\right)$.
@@
$\forall\lVert x\rVert_{2}\leq 1:\lVert J(x,\theta_{t})-J(x,\theta_{0})\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{\prime}C,$ (14)
$\lVert J(\theta_{t})-J(\theta_{0})\rVert_{2}\leq\frac{1}{\sqrt{n}}KC.$ (15)
We provide a proof in Appendix F.1. The main idea is to analyze $\frac{d}{dt}\lVert\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})\rVert_{2}$ and leverage the fact that the eigenvalues of the Hessian of the network scale as $O\left(\frac{1}{\sqrt{n}}\right)$. Consequently, these eigenvalues are dominated by the regularization term $\beta>0$, leading to the exponential decay. This is a considerable shift from prior proofs for the case $\beta=0$, where exponential decay is shown for $\lVert g(\theta_{t})\rVert_{2}$ (the training error) instead [^10] This implies exponential decay of $\lVert\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\rVert_{2}=\lVert J(\theta_{t})^{\top}g(\theta_{t})\rVert_{2}$..
Our result for $\beta>0$ does not require the minimum eigenvalue $\lambda_{\min}(\mathbf{\Theta})$ of the analytical NTK to be positive. For the unregularized case $\beta=0$ (and $\lambda_{\min}(\mathbf{\Theta})>0$), we recover the results of lee2019wide by an alternative proof. This highlights the importance of the gradient being in the row-span of the Jacobian $J(\theta_{t})$, ensuring that only the positive eigenvalues of $J(\theta_{t})^{\top}J(\theta_{t})$ contribute to the dynamics.
@@
$\sup_{t\geq 0}\lVert\theta_{t}-\theta^{\operatorname*{lin}}_{t}\rVert_{2}\leq C_{1}\frac{1}{\sqrt{n}},$ (16)
$\forall\lVert x\rVert_{2}\leq 1:\sup_{t\geq 0}\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t}^{\operatorname*{lin}})\rVert_{2}\leq C_{2}\frac{1}{\sqrt{n}}.$ (17)
See Appendix F.2 for a proof. Again, unlike previous proofs for $\beta=0$, which rely on the exponential decay of $\lVert g(\theta_{t})\rVert_{2}$, we present a more general and simpler approach that applies for $\beta\geq 0$. Specifically, we decompose $\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t}^{\operatorname*{lin}})\rVert_{2}$ into $\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t})\rVert_{2}$ and $\lVert f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t})-f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t}^{\operatorname*{lin}})\rVert_{2}$, and bound them using Theorem 4.2.

### v3

@@ Abstract
Performing gradient descent in a wide neural network is equivalent to computing the posterior mean of a Gaussian Process with the Neural Tangent Kernel (NTK-GP), for a specific prior mean and with zero observation noise. However, existing formulations have two limitations: (i) observation noise, since the NTK-GP assumes noiseless targets, leading to misspecification on noisy data; (ii) the equivalence does not extend to arbitrary prior means, which are essential for well-specified models. To address (i), we introduce a regularizer into the training objective, showing its correspondence to incorporating observation noise in the NTK-GP. To address (ii), we propose a shifted network that enables arbitrary prior means and allows obtaining the posterior mean with gradient descent on a single network, without ensembling or kernel inversion. We validate our results with experiments across datasets and architectures, showing that this approach removes key obstacles to the practical use of NTK-GP equivalence in applied Gaussian process modeling.
@@ 1 Introduction
The connection between wide neural networks and Gaussian Processes (GPs) via the Neural Tangent Kernel (NTK) (Jacot et al., 2018) provides a powerful framework for understanding the training dynamics of deep learning. The NTK describes how predictions of a neural network evolve under gradient-based optimization in the infinite-width limit and has primarily been used to analyze generalization, convergence, and expressivity of neural networks from a theoretical perspective. While this line of work has yielded insights into training dynamics, its practical utility for GP inference remains limited. In particular, the standard NTK-GP correspondence, as established by Lee et al. (2019), connects the output of a wide neural network trained with gradient flow on mean squared error (MSE) to the posterior mean of a GP with the NTK as its kernel — under restrictive conditions. Specifically, this result assumes (i) zero observation noise, leading to model misspecification when learning from real-world data, and (ii) a fixed prior mean given by a randomly initialized network, which is uninformative and not easily controllable. These constraints make the GP equivalence difficult to leverage in practice, as they limit the model’s ability to incorporate uncertainty and prior knowledge, two key concepts of principled Bayesian inference. In this work, we address these limitations and extend the theory in a way that enables practical posterior mean inference with NTK-GPs using a single neural network training run.
@@
Concerning the first limitation, observation — or aleatoric — noise is a fundamental component of Gaussian Processes and other probabilistic models (Williams and Rasmussen, 2006; Sáez de Ocáriz Borde et al., 2024; Ordoñez et al., 2024) that represents observation uncertainty. Measurements in data are subject to precision limits, and labels can carry errors due to annotation noise or ambiguities (Kendall and Gal, 2017). By capturing this variability, aleatoric noise enables models to produce predictions that reflect the randomness of the observed process (Williams and Barber, 1998; Bergna et al., 2024). In GPs, this is achieved through a variance term that accounts for label noise, contributing to robust and well-calibrated predictions.
@@
Hu et al. (2020) tried to address the problem of observation noise by introducing a regularizer penalizing the distance between network parameters at time $t$, $\theta_{t}$, and their initialization, $\theta_{0}$, demonstrating its efficacy in noisy settings. Their analysis, however, relied on the NTK framework and assumed that the network remains in a linear regime throughout training — a critical assumption that was left unproven. This oversight is significant, as the regularizer modifies the training dynamics and invalidates the application of previous results in Lee et al. (2019). Other recent works have used the regularizer proposed by Hu et al. (2020) to improve generalization and training stability. For instance, Nitanda and Suzuki (2020) employed it to constrain networks near their initialization under the NTK regime, while Suh et al. (2021) extended this to deeper networks for recovering ground-truth functions from noisy data. He et al. (2020) explored the regularizer’s role in Bayesian deep ensembles.
@@ 2 Preliminaries
The choice of parametrization in neural networks determines how signals propagate and how gradients scale with the network width. Let $\phi:\mathbb{R}\to\mathbb{R}$ denote the activation function [^1] We assume that $\phi$ and $\phi^{\prime}$ are Lipschitz.. In the standard parametrization, the layer outputs are defined (for $l=0,\ldots,L$) as:
@@
where $W^{l+1}\in\mathbb{R}^{n_{l+1}\times n_{l}}$ and $b^{l+1}\in\mathbb{R}^{n_{l+1}}$ are the weights and biases, respectively. The parameter vector $\theta\in\mathbb{R}^{p}$ is defined by stacking them. Here, $n_{0}:=d$ is the dimension of the input $x^{0}=x$, and $n_{L+1}:=k$ is the dimension of the output $f(x,\theta):=h^{L+1}$. For simplicity we only consider $k=1$. The initial weights $\theta_{0}$ are i.i.d. $W^{l}_{0,ij}\sim\mathcal{N}(0,\frac{\sigma_{w,l}^{2}}{n_{l}})$, and the biases as $b^{l}_{0,i}\sim\mathcal{N}(0,\sigma_{b,l}^{2})$ [^2] Note, that $W^{1}x^{0}$ sums over $d$ terms, and for $l\geq 1$, $W^{l+1}x^{l}$ sums over $n$ terms.. Under this parametrization, the norm of the Jacobian diverges for width $n_{l}\to\infty$. Without loss of generality, we will assume $n_{1}=\ldots=n_{L}=:n$.
@@
The NTK characterizes the evolution of wide neural network predictions as a linear model in function space. Given a neural network $f(x,\theta)$ parameterized by $\theta\in\mathbb{R}^{p}$, define the parameter-Jacobian in any $N$ points $\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\in\mathbb{R}^{d}$ as $J(\mathbf{x},\theta):=\frac{\partial f(\mathbf{x},\theta)}{\partial\theta}\in\mathbb{R}^{N\times p}$. Under NTK parametrization, the empirical NTK at two sets of inputs $\mathbf{x}^{\prime}_{1},\ldots,\mathbf{x}^{\prime}_{N^{\prime}}$ and $\mathbf{x}_{1},\ldots,\mathbf{x}_{N}$ is defined as:
$\hat{\mathbf{\Theta}}_{\mathbf{x}^{\prime},\mathbf{x}}:=J(\mathbf{x}^{\prime},\theta_{0})J(\mathbf{x},\theta_{0})^{\top}\in\mathbb{R}^{N^{\prime}\times N}.$ (3)
The empirical NTK depends on the randomly initialized $\theta_{0}$. Interestingly, as shown by Jacot et al. (2018), in the infinite width limit, the empirical NTK converges to a deterministic (i.e., independent of $\theta_{0}$) kernel $\mathbf{\Theta}$ (the analytical NTK) at initialization and remains constant during training (with unregularized gradient flow). A GP defined by the analytical NTK is referred to as the NTK-GP.
@@ 3 Motivation
Our work builds on a well-known result (Lee et al., 2019) which shows that, for any sufficiently [^3] In our proofs, we make the notion of ‘sufficiently wide’ formal. wide neural network initialized to parameters $\theta_{0}$, performing gradient descent on the loss
$\mathcal{L}^{0}(\theta):=\frac{1}{2}\sum_{i=1}^{N}(f(\mathbf{x}_{i},\theta)-\mathbf{y}_{i})^{2}$ (4)
@@
We answer this question in the affirmative. Specifically, Section 4 addresses limitation (i), observation noise, by showing that adding a regularizer to the training objective induces aleatoric noise $\beta$ in the NTK-GP posterior mean (Theorems 4.1 and 4.3), while Section 5 addresses limitation (ii), arbitrary prior means, via a shifted network construction that enables user-specified priors (Theorem 5.1).
@@ 4 Observation Noise through Regularized Gradient Descent
The introduction of weight-space regularization not only impacts the generalization properties of wide neural networks (Hu et al., 2020), but also fundamentally alters their training dynamics. In the results proposed by Lee et al. (2019) —where the unregularized gradient flow is studied— these dynamics are crucial to understanding how networks converge to solutions that (in the mean posterior sense) align with Bayesian inference. This section addresses limitation (i), observation noise, by analyzing the training dynamics of wide neural networks under regularized gradient flow, with a particular focus on how regularization affects network linearization during training. We demonstrate that the training trajectory stays arbitrarily close to its linearized counterpart, providing both theoretical justification and insights into how regularization modifies the gradient flow.
Define $f(\theta):=f(\mathbf{x},\theta)$, $g(\theta):=f(\mathbf{x},\theta)-\mathbf{y}\in\mathbb{R}^{N}$, $J(\theta):=J(\mathbf{x},\theta)\in\mathbb{R}^{N\times p}$ in the training points [^4] We assume that $\mathbf{x}_{i}\neq\mathbf{x}_{j}$ for $i\neq j$.. We will consider (for NTK parametrization [^5] See Appendix E.2 for standard parametrization.) the regularized training loss
@@
We derive this closed-form solution (for gradient flow and gradient descent) in Appendix D. Note that the statements are high-probability results with respect to the random initialization: for any $\delta_{0}>0$, the probability can be made at least $1-\delta_{0}$ by choosing the width $n$ sufficiently large. The required scaling of $n$ with $\delta_{0}$ is made precise in Appendices A and F.
Moreover, if we consider the network initialization as a random variable, the trained linearized network converges to a normal distribution for layer width $n\to\infty$. Its mean corresponds to the posterior mean of the NTK-GP with prior mean 0. However, similarly to Lee et al. (2019); He et al. (2020), the covariance cannot be interpreted as the posterior covariance of an NTK-GP. See Appendix I for the exact formulas.
In the remainder of this section, we show that, with high probability over initialization, the training dynamics of the network (at any test point $\mathbf{x^{\prime}}$) closely follow the output of the linearized network for large enough hidden layer width $n$. First, we leverage the fact that the Jacobian is locally bounded and locally Lipschitz, with a Lipschitz constant of $O\left(\frac{(\log n)^{c}}{\sqrt{n}}\right)$ (see Appendix A for details). Next, we prove that the distance between the parameters and their initialization $\lVert\theta_{t}-\theta_{0}\rVert_{2}$ is O(1)O(1) [^8] On average, this implies that the distance of individual parameters barely changes.. This step is particularly challenging because proofs for $\beta=0$ (in Eq. (6)) rely on the MSE $\lVert g(\theta_{t})\rVert_{2}$ decaying exponentially to 0. As this is not the case for $\beta>0$, we instead show that $\lVert\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})\rVert_{2}$ decays to 0 exponentially fast.
@@ 4.1 Exponential Decay of the Regularized Gradient and Parameter Proximity to Initialization
Building upon the established Lipschitz continuity of the Jacobian (in Appendix A), we analyze the effect of regularized gradient flow on the norm of the gradient and the parameters’ closeness to their initialization. We begin by demonstrating that the norm of the gradient of the regularized loss decays exponentially over time. This result directly implies that the distance between the parameters and their initialization is bounded by a constant, i.e., $\lVert\theta_{t}-\theta_{0}\rVert_{2}=O(1)$ [^9] The bounding constant is independent of the outcome in the high probability event for which this holds.. Furthermore, leveraging the Lipschitzness of the Jacobian established in Lemma A.1, we show that the Jacobian remains close to its initial value, with deviations scaling as $O\left(\frac{(\log n)^{c}}{\sqrt{n}}\right)$.
@@
$\forall\lVert x\rVert_{2}\leq 1:\lVert J(x,\theta_{t})-J(x,\theta_{0})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}C,$ (14)
$\lVert J(\theta_{t})-J(\theta_{0})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}KC.$ (15)
We provide a proof in Appendix F.1. The main idea is to analyze $\frac{d}{dt}\lVert\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})\rVert_{2}$ and leverage the fact that the eigenvalues of the Hessian of the network scale as $O\left(\frac{(\log n)^{c}}{\sqrt{n}}\right)$. Consequently, these eigenvalues are dominated by the regularization term $\beta>0$, leading to the exponential decay. This is a substantial shift from prior proofs for the case $\beta=0$, where exponential decay is shown for $\lVert g(\theta_{t})\rVert_{2}$ (the training error) instead [^10] This implies exponential decay of $\lVert\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\rVert_{2}=\lVert J(\theta_{t})^{\top}g(\theta_{t})\rVert_{2}$..
Our result for $\beta>0$ does not require the minimum eigenvalue $\lambda_{\min}(\mathbf{\Theta})$ of the analytical NTK to be positive. For the unregularized case $\beta=0$ (and $\lambda_{\min}(\mathbf{\Theta})>0$), we recover the results of Lee et al. (2019) by an alternative proof. This highlights the importance of the gradient being in the row-span of the Jacobian $J(\theta_{t})$, ensuring that only the positive eigenvalues of $J(\theta_{t})^{\top}J(\theta_{t})$ contribute to the dynamics.
@@
$\sup_{t\geq 0}\lVert\theta_{t}-\theta^{\operatorname*{lin}}_{t}\rVert_{2}\leq C_{1}\frac{(\log n)^{c}}{\sqrt{n}},$ (16)
$\forall\lVert x\rVert_{2}\leq 1:\sup_{t\geq 0}\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t}^{\operatorname*{lin}})\rVert_{2}\leq C_{2}\frac{(\log n)^{c}}{\sqrt{n}}.$ (17)
Where $\theta^{\operatorname*{lin}}_{t}$ are the parameters at time $t$ of the linearized network $f^{\operatorname*{lin}}_{\theta_{0}}$. See Appendix F.2 for a proof. Again, unlike previous proofs for $\beta=0$, which rely on the exponential decay of $\lVert g(\theta_{t})\rVert_{2}$, we present a more general and simpler approach that applies for $\beta\geq 0$. Specifically, we decompose $\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t}^{\operatorname*{lin}})\rVert_{2}$ into $\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t})\rVert_{2}$ and $\lVert f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t})-f^{\operatorname*{lin}}_{\theta_{0}}(x,\theta_{t}^{\operatorname*{lin}})\rVert_{2}$, and bound them using Theorem 4.2.